What is the horizontal range of projectile if the object is fired from a building 400m high and has an initial velocity of 350m/s? The angle of elevation of the projectile is 250?
1
Expert's answer
2021-08-17T16:16:02-0400

"θ=25° \\
R= u \cos θ×T\\
T=total \space time \space of \space flight\\
T=\frac{2u sin θ}{R}\\
R=u \cos θ× \frac{2u sin θ}{g} \\
R =\frac{u^2 sin 2θ}{g}\\
R =\frac{350^2 sin 50}{9.8}=9575.55 m"

This R is the horizontal range of the projectile when it is at the level of  400 m above the ground, 

At this point  the velocities are again u and making an angle 25° with the horizontal but pointing downwards

"400=350 sin 25×t+\frac{1}{2}×9.8×t^2\\
t=2.497 s"

During this time interval, the extra range covered will be

"r=u cos θ×t \\
=350 cos 25×2.497 \\
=792.067 m"

total range (horizontal)"=r+R=10367.61 m=10.367 km"